### v1

# An ultra-broadband photonic-chip-based traveling-wave parametric amplifier

An ultra-broadband photonic-chip-based traveling-wave parametric amplifier
Nikolai Kuznetsov Institute of Physics, Swiss Federal Institute of Technology Lausanne (EPFL), CH-1015 Lausanne, Switzerland Center of Quantum Science and Engineering (EPFL), CH-1015 Lausanne, Switzerland    Alberto Nardi IBM Research Europe, Zurich, Säumerstrasse 4, CH-8803 Rüschlikon, Switzerland Institute of Physics, Swiss Federal Institute of Technology Lausanne (EPFL), CH-1015 Lausanne, Switzerland    Alisa Davydova Institute of Physics, Swiss Federal Institute of Technology Lausanne (EPFL), CH-1015 Lausanne, Switzerland Center of Quantum Science and Engineering (EPFL), CH-1015 Lausanne, Switzerland    Mikhail Churaev Institute of Physics, Swiss Federal Institute of Technology Lausanne (EPFL), CH-1015 Lausanne, Switzerland Center of Quantum Science and Engineering (EPFL), CH-1015 Lausanne, Switzerland    Johann Riemensberger johann.riemensberger@ntnu.no Institute of Physics, Swiss Federal Institute of Technology Lausanne (EPFL), CH-1015 Lausanne, Switzerland Center of Quantum Science and Engineering (EPFL), CH-1015 Lausanne, Switzerland Department of Electronic Systems, Norwegian University of Science and Technology, 7491 Trondheim, Norway    Paul Seidler pfs@zurich.ibm.com IBM Research Europe, Zurich, Säumerstrasse 4, CH-8803 Rüschlikon, Switzerland    Tobias J. Kippenberg tobias.kippenberg@epfl.ch Institute of Physics, Swiss Federal Institute of Technology Lausanne (EPFL), CH-1015 Lausanne, Switzerland Center of Quantum Science and Engineering (EPFL), CH-1015 Lausanne, Switzerland
Optical amplification, crucial for modern communication, primarily relies on erbium-doped fiber amplifiers (EDFAs) to enhance signals without distortion. However, EDFAs only cover a portion of the low-loss spectrum of optical fibers. Optical parametric amplification, utilizing material nonlinearities, offers expanded capacity, but has faced limitations in gain and bandwidth. Addressing this, we present a high-gain continuous-traveling-wave parametric amplifier, surpassing the erbium gain window. Leveraging low-loss gallium phosphide-on-insulator photonic integrated waveguides, we achieve up to 35 dB of parametric gain, with a fiber-to-fiber net gain exceeding 10 dB across a wide wavelength range of approximately 140 nm centered at 1550 nm. We demonstrate the capability to amplify ultra-weak signals, extending over 6 orders of magnitude of input powers, optical frequency combs, and communication signals, highlighting the practical potential of our photonic-chip-based amplifier.
Introduction. Optical fibers and the broadband amplification of traveling- and continuous-wave (CW) optical signals have provided pivotal advancements in modern science and technology, particularly in the domain of optical communications for the transmission of information over long distances. A crucial breakthrough for intercontinental optical fiber networks was the introduction of erbium-doped fiber amplifiers (EDFAs) [1, 2], which eliminated the need for frequent electronic signal regeneration and played a pivotal role in the expansion of the World Wide Web. Efforts have been made to extend the bandwidth of the optical gain by using Raman-assisted EDFAs [3] and semiconductor optical amplifiers (SOAs) [4]. In contrast to amplifiers based on rare-earth-doped fibers or the more recently developed rare-earth-doped photonic integrated circuits (PICs) [5, 6], optical parametric amplifiers (OPAs) rely on the intrinsic nonlinearities of optical materials to generate gain. Such traveling-wave parametric amplifiers (TWPAs) were originally developed in the microwave domain, where they, e.g., provide quantum-limited amplification for qubit readout [7].
OPAs have unique properties that set them apart from other means of amplification [8, 9]. The shape of the amplification window is entirely determined by the dispersion and is limited only by the transparency of the material. The central frequency is adjustable, offering flexibility in signal processing. Parametric amplification provides the capability of frequency conversion through the creation of a concomitant idler wave carrying the same information as the input signal but at a different frequency [10, 11]. The reversal of the optical phase of the idler relative to that of the signal can be utilized to compensate for dispersion and mitigate nonlinear effects in transmission systems [12, 13, 14]. Kerr nonlinearity offers a nearly instantaneous response, allowing rapid operation of the OPA. In contrast to the gain from erbium, parametric gain can be adjusted by varying the optical drive power while maintaining a low noise figure, a behavior essential for amplification of weak input signals. Moreover, optical TWPAs operate near the quantum noise limit and offer the ability to perform noiseless phase-sensitive amplification and thus have the potential to increase the span length of long-haul fiber optical links [15, 16]. Lastly, OPAs are unidirectional, making them resistant to optical feedback and parasitic lasing, reducing the need for optical isolators.
To date, time-continuous operation of an OPA with substantial gain has only been demonstrated in fiber-based or bulk-crystal systems, for example, using hundreds of meters of highly nonlinear fibers [17, 18, 8, 19, 20] that require mitigation of strong detrimental Brillouin scattering, or using mechanically cut lithium niobate waveguides [21, 22], for which the $\chi^{(2)}$ nonlinearity requires periodic poling to generate quasi-phase matching and the performance is limited by the available wafer size, as the mode area is large and the bending losses are excessive. In contrast, photonic integrated circuits (PICs) offer the possibility of a significantly increased parametric gain coefficient $g=\sqrt{(\gamma P_{\textrm{P}})^{2}-(\kappa/2)^{2}}$ [8, 23], which is determined by the effective nonlinearity $\gamma={n_{2}\omega_{\textrm{P}}}/{cA_{\mathrm{eff}}}$ and the phase-mismatch parameter $\kappa$; here, $n_{2}$ is the Kerr nonlinearity, $A_{\mathrm{eff}}$ is the effective mode area, $\omega_{\textrm{P}}$ is the pump angular frequency, and $P_{\textrm{P}}$ is the pump power. With the proper choice of material, PICs can exhibit a strong Kerr nonlinearity while simultaneously featuring a high linear refractive index that substantially reduces the mode area. The material should also have negligible two-photon absorption and a high optical-damage threshold. Integrated photonic platforms that have been studied so far include silicon [24, 25, 26, 27, 28], chalcogenides [29], Si7N3 [30], AlGaAs [31], and highly doped silica [32]. None of these materials have yet produced net gain in the continuous pumping regime, as required for time-varying signals used in most applications. Recently, triggered by the advances in ultra-low-loss meter-long Si3N4 integrated photonic waveguides [33], it was possible to achieve CW parametric amplification with net gain [34, 35]. However, this approach still requires coping with high pump-power levels, meter-long waveguides, limited gain, and narrow bandwidth. Here, we address these challenges by demonstrating a compact, ultra-broadband, high-gain, PIC-based optical TWPA.
Results. We use thin-film gallium phosphide (GaP) on silicon dioxide to create a TWPA comprising a dispersion engineered waveguide operating at a pump wavelength near 1550 nm (see the Supplementary Material for details of the fabrication). GaP combines a high optical refractive index $(n=3.05)$ with a strong Kerr nonlinearity ($n_{2}=1.1\times 10^{-17}$ m2 W-1) and an indirect bandgap ($E_{\mathrm{g}}=2.24$ eV)  [36] sufficiently large to mitigate two-photon absorption at telecommunication wavelengths. This exceptional confluence of properties has facilitated the generation of low-threshold frequency combs [37] and dissipative Kerr solitons [38]. We estimate the effective nonlinearity of the GaP waveguides in this work to be $\gamma=165$ W-1m-1, which is more than 300 times larger than the value of 0.51 W-1m-1 reported for Si3N4 waveguides [35], facilitating a 35-fold reduction in waveguide length and a 60-fold reduction in device footprint. Advances in fabrication techniques [37, 38] allow us to reduce GaP waveguide optical propagation losses below 1 dB cm-1 and achieve a high bonding yield and low defect density. These developments are pivotal for successful fabrication of cm-long waveguide spirals with a footprint of only 500 $\times$ 500 $\mu$m2 (Figs. 1(C - G) & Supplementary Material).
We operate the TWPA with a single pump to amplify signals via degenerate four-wave-mixing (FWM), exploiting the optical Kerr effect (Fig. 1(a)). The optical energy is redistributed by annihilating two pump photons to amplify signal photons at a separate frequency, while simultaneously generating a phase-conjugated idler at a frequency equidistant from the pump laser but offset in the opposite direction. The near instantaneous action of the optical Kerr effect necessitates phase-matching in the waveguide for efficient amplification [8], i.e., $\kappa=\Delta\beta+2\gamma P_{\textrm{P}}$ is small, where $\Delta\beta$ is the linear propagation mismatch as given by $\Delta\beta\approx\beta_{2}\left(\omega_{\textrm{S}}-\omega_{\textrm{P}}\right)^{2}+\beta_{4}\left(\omega_{\textrm{S}}-\omega_{\textrm{P}}\right)^{4}/12+...$, $\beta_{2}$ and $\beta_{4}$ are the second- and fourth-order derivatives with respect to the angular frequency $\omega$ of the optical propagation constant $\beta$ evaluated at the pump frequency $\omega_{\textrm{P}}$, and $\omega_{\textrm{S}}$ is the signal angular frequency. Only even order dispersion terms contribute to the shape of the amplification spectrum. The largest optical bandwidth is achieved when the dispersion is anomalous ($\beta_{2}<0$) and the linear phase mismatch $\Delta\beta$ is compensated by the nonlinear phase mismatch $2\gamma P_{\textrm{P}}$ originating from self- and cross-phase modulation. The peak gain in a waveguide of length $L$ is defined by $G_{\textrm{S}}=1+\left[\sinh(-\Delta\beta L_{\textrm{eff}}/2)\right]^{2}$, where $L_{\textrm{eff}}=(1-\exp(-\alpha L))/\alpha$ is the effective length and $\alpha$ is the linear propagation loss. Fig. 1(b) shows a comparison of the gain spectrum calculated for an optimized GaP optical TWPA with that of C-band and L-band EDFAs. The optimized GaP waveguide has a cross-section of $789\times 299$ $\text{nm}^{2}$, for which $\beta_{2}=-16$ fs2mm-1 and $\beta_{4}=3547$ fs4mm-1. The region of signal and idler gain extends from around 1300 nm in the optical O-band to 1900 nm, well beyond the longest wavelengths used in silica fiber communications. We find that the waveguides used in this work exhibit a stronger anomalous dispersion than expected from our simulations (see Supplementary Material).
Figure 1: Optical continuous-wave parametric amplification in a integrated GaP photonic waveguide. (a) Principle of broadband optical parametric amplification in an integrated spiral waveguide. Strong pump and small signal waves co-propagate through the waveguide, leading to the amplification of the signal and generation of the third wave, the idler. The frequency of the signal wave can be chosen over a wide range within the amplification bandwidth. (b) Comparison of the typical amplification bandwidth of state-of-the-art EDFAs (green: C-band, yellow: L-band) with the amplification bandwidth of integrated GaP parametric amplifiers (red: achieved in this work, blue: if further optimized). (c) Optical microscope image of the fabricated photonic chip with multiple spiral waveguides and other test structures. (d) Enlarged optical microscope image of a 5-cm-long spiral waveguide. Accounting for the chip width of 0.55 cm, the total waveguide length is 5.55 cm. (e) Scanning electron microscope image of a few waveguide turns. (f) Scanning electron microscope image of an inverse taper facet, similar to the one used in this work, used to couple light into a GaP waveguide. (g) Material stack of the fabricated chip.
We measure the CW amplification spectrum of a 5.55 cm-long GaP spiral waveguide with a height of 299 nm and a design width of 780 nm using the same methods as described in Ref. [35]. Figs. 2(A, B) depict the experimental setup and recording scheme used to measure the parametric gain (see Supplementary Material). Fig. 2(c) shows the power-calibrated amplification spectrum. In the spectral region around the pump wavelength, we observed a broadened peak due to the leakage of the amplified spontaneous emission from the EDFA installed in the optical path of the pump; our tunable filter has a bandwidth of 1 nm and a finite suppression ratio of 30 dB. The frequency conversion and amplification spectra (Figs. 2(D, E)) are determined by comparing the output spectra recorded as the the signal laser is swept at various pump power levels (red to orange) with the output spectrum obtained when the pump is switched off (black). Increasing the input power from 1.00 W to 4.13 W, we observe increasing gain, conversion efficiency, and amplification bandwidth, following established theory [39, 8]. The generated idler spectrum spans from 1550 nm to approximately 1480 nm at the highest pump power. In Fig. 2(c), spurious light generated outside the scan range of the signal laser (outer shaded regions) originates from non-degenerate FWM. Hence, we only calculate the internal idler conversion efficiency (Fig. 2(d)) from 1480 nm to 1540 nm and the internal signal gain from 1560 nm to 1630 nm (Fig. 2(e)). We separately measure the passive transmission loss of the TWPA, i.e., the insertion loss, by bypassing the lensed fibers and the chip; the ratio of the spectrum obtained with the bypass to the spectrum transmitted through the chip is plotted as a black line in Fig. 2(e). To calculate the off-chip net gain, we subtract this 10 – 12 dB of insertion loss, revealing net-gain of up to $G_{\textrm{S}}=25$ dB, with $G_{\textrm{S}}>10$ dB over a bandwidth of 70 nm. Assuming a symmetric gain spectrum on the short-wavelength side of the pump, we estimate the 10 dB off-chip gain region to exceed 140 nm, or 17 THz. This bandwidth is almost ten times larger than reported in prior work on Si3N4 [35], and nearly three times larger than the bandwidth of a single C-band EDFA.
Figure 2: Observation of a broadband continuous wave parametric amplification in a GaP spiral waveguide. (a) Schematic of the experimental setup. EDFA: erbium-doped fiber amplifier, (d)VOA: (digital) variable optical attenuator, PC: polarization controller, PM: power meter, OSA: optical spectrum analyzer. (b) Principle of the "Max Hold" setting of the OSA used in broadband gain measurements. With every new scan, the OSA records and updates only the highest values across the measurement span, while the wavelength of the signal laser is slowly swept to cover the whole amplification bandwidth. (c) Measured amplification spectra for different off-chip pump powers. The resolution bandwidth of the OSA is 2 nm. EDFA and DVOA in the signal path are not used in this experiment. (d, e) Calculated idler conversion efficiency (CE) and signal amplification spectra for fiber-coupled input pump powers up to 4.13 W, respectively. (f) Off-chip signal power at wavelength 1605 nm at the output of the amplifier as a function of signal and pump powers at the input fiber. The dashed line shows the boundary when the input signal power is equal to the output signal power. (g) Measurement of the optical TWPA noise figure as a function of signal input power. The dashed line indicates the 3 dB quantum limit of a phase-insensitive amplifier.
To investigate the maximum signal output power and the saturation power of the amplifier, we perform a power sweep of the signal at a wavelength of 1605 nm (Fig. 2(f)). Here we use a different spiral with a designed waveguide width of 790 nm; this spiral is used also in the remaining experiments. The measurements are carried out with the resolution bandwidth of the OSA set at 0.1 nm. At a pump power of 4.43 W, the saturated off-chip signal output power exceeds 125 mW, corresponding to a power of 220 mW on chip at the waveguide output. Hence, the on-chip power conversion efficiency reaches 9 %, which is limited by the waveguide propagation losses. Here, we also demonstrate the ability to achieve low-noise amplification with high gain over a large dynamic range. The linear, small-signal regime of amplification extends up to a signal input power of almost 0 dBm at the highest pump power level of 4.43W. At 2.01 W pump power, the amplification is linear over the entire range of input powers measured, extending over six orders of magnitude and limited only by the capabilities of our measurement setup; in principle, parametric amplification poses no lower limit on the signal input power. We calculate the amplifier noise figure from the measured signal amplification and the spontaneous parametric fluorescence background (see Supplementary Material) [40]. The off-chip noise figure approaches 6 dB in the small-signal gain regime. Since degradation of the signal-to-noise ratio (SNR) at the input facet is irreversible, we take into account the coupling losses of 2.5 dB at the chip input facet and find the on-chip noise figure of less than 4 dB for a wide range of signal powers below saturation (Fig. 2(g)). The degradation of the noise figure at higher signal input powers is attributed to gain saturation, in other words, the signal gain becomes weaker while the parametric fluorescence is still within the small-signal regime and undergoes a higher amplification.
Figure 3: Amplification of optical frequency combs using a GaP optical TWPA. (a) Schematic of amplification and frequency conversion of EO frequency comb (top) and soliton microcomb (bottom). (b) Amplified EO comb spectra (red) and fiber-coupled input spectrum of OPA (blue) showing more than 20 dB fiber-to-fiber net gain. Envelopes are added for better visibility. (c) Same as (b) but for amplification of 100 GHz microcomb produced by single soliton state in a Si3N4 microring cavity. The inset shows the appearance of frequency-converted idler comb lines as a result of DFWM. Increased noise floor at longer wavelengths on both spectra is due to the proximity to the end of the OSA measurement range. The resolution bandwidth of the OSA is 0.02 nm in both experiments.
To highlight the application potential of such a broadband optical TWPA, we perform additional amplification experiments using low-power and high-repetition-rate frequency comb sources as signals. In Fig. 3(a), we schematically show amplification and idler generation for two types of injected frequency combs, a narrowband electro-optic (EO) comb [38] and a broadband dissipative Kerr soliton comb [41]. The EO comb is centered around 1600 nm and has a 16 GHz line spacing and a total signal power of 10.5 $\mu$W. When transmitted through the amplifier together with a 4.12 W pump at 1550 nm, we observe more than 20 dB fiber-to-fiber net gain (Fig. 3(b)). The reference dataset (blue) is measured by bypassing the lensed fibers and the photonic chip. Note that, due to the nearly instantaneous nature of parametric amplification, we do not use a dispersive fiber compression stage during the EO comb preparation before the amplifier to avoid high signal peak powers and amplifier saturation. We perform the same experiment with the frequency comb formed by a dissipative Kerr soliton state [41] generated in a Si3N4 microring resonator. The total power of the soliton comb is 76 $\mu$W, and the repetition rate is 100 GHz. In the spectral region of the highest gain, far from the soliton pump located at 1552.8 nm, the input power of individual lines decreases to the level of a few nW, while lines with the highest initial power, located closer to the pump, experience smaller gain. As expected, an idler comb mirrored with respect to the pump frequency is generated, and the line density of the output spectrum is doubled (Fig. 3(c), red). Additional lines can be observed throughout the entire spectrum and, most notably, closer to the pump; we attribute their appearance to numerous FWM processes between various interacting lines. Additional lines that are located at frequencies far outside the bandwidth of the input comb arise due to cascaded non-degenerate FWM processes. A small feature in the amplified spectra around 1651 nm is observed with all available waveguides and can be attributed to Raman scattering in GaP [37, 42]. The flattened spectra resulting from amplification provide clear evidence that the GaP OPA can handle the simultaneous input of a few hundred lines over a broad bandwidth, and the parametric gain remains as high as for a single input frequency if operated in the small-signal regime. Moreover, we again observe amplification of very small input signals, with the lowest comb lines being as weak as 1 nW, something that is not possible with typical EDFAs. The amplification of such small signals may be especially attractive for applications such as LiDAR [43] or free-space communications.
Figure 4: Coherent data transmission experiment. (a) Simplified schematic of the experimental setup used in the communication experiment AWG: arbitrary waveform generator. FBG: fiber Bragg grating. WDM: wavelength-division multiplexer. OSC: oscilloscope. (b) Calibrated optical spectra measured at the output of the chip using the same OSA as described above (not shown here in (a)); resolution bandwidth is set to 0.2 nm for a faster spectra acquisition. (c, d, e) Constellation diagrams and modulation error ratios (MERs) for three measurements, respectively: fiber-to-fiber reference, amplified signal, generated idler.
To further demonstrate the potential of the GaP OPA for use in coherent communication, we set up a 10 GBd communication line, as shown in Fig. 4(a). We send a QPSK-encoded pseudo-random bit sequence at a signal wavelength of 1605 nm (L-band) to our OPA and then perform heterodyne measurements of the amplified signal. Additionally, using another laser as a local oscillator, we measure modulation transfer to the idler located at approximately 1502 nm (S-band). The signal wavelength is chosen to be in the region of maximum gain, while also considering the limitations of the available equipment. Fig. 4(b) shows corrected OPA output spectra with the pump switched on and off. The internal gain achieved exceeds 30 dB, which is enough to compensate for coupling and propagation losses on the chip (which barely exceed 10 dB in total at the selected signal wavelength) and even for 10 dB of losses on the splitter before the chip and 4 dB of losses in the filtering section and wavelength-division multiplexer, so we achieve a positive net gain accounting for all losses present in the system (see Supplementary Material). The received signal power is 26 $\mu$W and the idler power is 32 $\mu$W; the idler power is higher than the signal power due to lower losses at the idler wavelength. As a reference, we measure the signal directly sent from the transmitter to the receiver through a fiber, bypassing all the other components without attenuation. We digitally post-process and analyze the data and plot constellation diagrams for all three measurements – for the fiber reference, signal, and idler (Figs. 4(C, D, E), respectively). As a figure of merit, for each measurement we estimate the modulation error ratio (MER) [44], defined as the sum of the squares of the magnitude of the ideal symbol vectors divided by the sum of the squares of the magnitude of the symbol error vectors. The idler conversion efficiency is high enough in the GaP OPA to generate an idler of the same power level as that of the amplified signal. This eliminates any need for additional idler post-amplification, as it is sufficiently strong to be used directly for inter-band signal translation.
Discussion. In summary, using an integrated GaP photonic platform, we achieve both high net gain and broadband operation of an OPA. The fiber-to-fiber net gain exceeds 20 dB, and the combined signal and idler gain bandwidth is 140 nm, significantly greater than the bandwidth of EDFAs. We demonstrate the practical application potential of GaP waveguides for the amplification of frequency combs, optical coherent data streams, and weak optical signals over more than 60 dB of dynamic range. Further reducing the optical propagation losses will greatly lower the required pump power and may even allow direct pumping of the OPA with semiconductor lasers.
Our on-chip GaP OPA solves some of the most pivotal challenges that have prevented the widespread adoption of fiber OPA systems. The dispersion of the waveguides can be designed as needed because it is defined lithographically, and the waveguide’s short length avoids a reduction in bandwidth [45] and reduces sensitivity to fabrication imperfections, in contrast to optical fibers that are hundreds of meters long. The low Brillouin gain of integrated waveguides eliminates the technical complexity of adding additional pump phase modulators [19, 20, 15]. In principle, the pump frequency can be freely chosen over a wide frequency range, as it is not restricted by an excitation spectrum as in doped amplifiers. Also, in contrast to existing EDFA technology, the optical TWPA is unidirectional, making it more robust against parasitic reflections. Lastly, the TWPA may achieve a lower noise figure, which can be reduced below the quantum limit if operated in the phase-sensitive regime [16, 34].
Our results signal the emergence of practical optical TWPAs with large bandwidth, high gain, and small footprint that could eventually be employed in future optical communication systems or other applications. Broadband optical TWPAs may be ideal candidates to drive next-generation fiber systems relying on doubly antiresonant hollow-core fibers that offer much increased spectral bandwidth for optical communications compared to silica fibers. Nonlinear distortions in parametrically amplified fiber links may also be mitigated by alternately propagating the signal and phase-conjugated idler through successive fiber link sections [46]. In general, the ability to amplify weak optical signals over a large dynamic range may be decisive for a variety of applications, including coherent communications, LiDAR and sensing.
Author contributions: N.K. and J.R. performed numerical simulations of dispersion and parametric gain, and designed spiral waveguides. A.N. fabricated the sample. J.R. and N.K. performed the initial linear characterization of the devices. N.K. and A.N. carried out the broadband gain measurements. N.K. and A.D. conducted frequency comb amplification experiments. N.K. and M.C. executed the communication experiment. N.K. carried out the rest of the measurements presented in this work. N.K. and J.R. prepared the manuscript with contributions from all authors. J.R., P.S. and T.J.K supervised the work.
Funding information: This work was supported by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 812818 (MICROCOMB), by the Swiss National Science Foundation (SNF) under grant number 192293, and by the Air Force Office of Scientific Research under award number FA9550-19-1-0250.
Acknowledgments: The sample was fabricated at the Binnig and Rohrer Nanotechnology Center (BRNC) at IBM Research Europe, Zurich.
Disclosures: J.R., N.K. and T.J.K. are inventors on European Patent Application No. 24165690.9 submitted by Swiss Federal Institute of Technology Lausanne (EPFL) that covers “Optical parametric amplifier apparatus and optical signal transmission apparatus, and applications thereof”. Other authors declare no conflicts of interest.
Data availability: All experimental datasets and scripts used to produce the plots in this work will be uploaded to Zenodo repository upon publication of this preprint.